A dryer, working continuously, is being fed with a wet material that has 60% solids and 40% water. 1000 kg / h of product with 10% humidity come out of the dryer. The dryer receives dry air (79% mol of N2 and 21% mol of oxygen) through a 70 cm internal diameter duct. Air enters at a speed of 60 km / h at a pressure of 585 mmHg of absolute pressure and 100 ° C.

Calculate:
a) The amount of wet material that enters the drying in Kg / h.
b) Evaporated water in Kg / h.
c) The amount of humid air that comes out of the dryer in Kg / h.
d) Air humidity in Kg of water / Kg of dry air
Answer #1

Given that,

A dryer, working continuously, is being fed with a wet material that has 60% solids and 40% water.

Let the wet feed be F

x_{F} be the fraction of feed that is water x_{F}=0.4

The amount of product is given as P=1000kg/h which has a humidity of 10%

x_{P} be the fraction of product that is water x_{P}=0.1

The dryer has an internal diameter of D=70cm=0.7m

The dryer recieves dry air.

The air enters at a velocity of v=60km/h=16.67m/s

and at a pressure of p=585mmHg=77993.6Pa and temperature of T=100^{0}C=373K

Now we will find the solutions of the questions,

Doing balance on the dry solid coming inside and going outside we get,

(1-x_{F})F=(1-x_{p})P

\Rightarrow (1-0.4)F=(1-0.1)*1000

\Rightarrow F=1500kg/h

So

amount of wet material that enters the dryer is 1500kg/h

b).Taking the overall mass balance on the dryer we get,

wet \: material=water\: evaporated+dried\: product

So,

1500=water\: evaporated+1000

\Rightarrow water\: evaporated=500kg/h

So, the water evaporated will be at the rate of 500kg/h

c)

The volumetric flow rate of the air entering can be calculated as,

q=vA

Where A is the area of the cross section of the dryer, A=\frac{\pi }{4}D^{2}=\frac{\pi }{4}0.7^{2}=0.385m^{2}

So,

q=vA=16.67*0.385=6.414m^{3}/s=23090.75\: m^{3}/h

We should multiply the volumetric flow rate with it's density to find the mass flow rate of the dry air entering,

The average molecular weight of air can be calculated as,

M=0.79*M_{N_{2}}+0.21*M_{O_{2}}=0.79*28+0.21*32=28.84g/mol=28.84*10^{-3}kg/mol

Considering the air to be ideal gas, density is given as,

\rho =\frac{pM}{RT}=\frac{77993.6*28.84*10^{-3}}{8.314*373}=0.725kg/m^{3}

So, mass flow rate is,

m=q*\rho=23090.75*0.725=16740.8\: kg/h

Now we have to calculate the flow rate of humid air that comes out of the dryer which is,

humid \: air=water\: evaporated+dry\: air

\Rightarrow humid \: air=500+16740.8=17240.8kg/h

d) So now the humidity will be,

Y=\frac{water}{dry\: air}=\frac{500}{16740.8}=0.02987\; kg\; water/kg\; dry\; air